steel pipe 150mm external diameter conveys steam at a temperature of 260 degrees and is covered by two layers of lagging, each 50mm thick. The thermal conductivity coefficient of the inside layer of lagging is 0.086W/m.K while that of the outside layer is 0.055W/m.K. The outside surface temperature of the steel pipe can be taken as the same as the temperature of the system. The ambient temperature is 27 degrees. Calculate the heat lost per hour for pipe length 30m. Calculate the interface temperature between the two layers of lagging


1
Expert's answer
2021-11-29T00:18:18-0500

 heat lost


"Q = \frac{\triangle T}{R_{th}}=\frac{260-27}{0.086+0.055} = 1652.48W=1.65kW"